# Discrete conformal structures on surfaces with boundary (II)—Rigidity and Existence

### v1

@@ 6.3. Existence of the mixed discrete conformal structure $\mathrm{II}$
$\widetilde{L}_{1}\rightarrow$ $\frac{c_{k}e^{f_{i}+f_{j}}}{\sqrt{(-1+\xi_{jk}e^{f_{j}+f_{k}})^{2}-1}\cdot\sqrt{(1+\xi_{ik}e^{f_{i}+f_{k}})^{2}-1}}$
$=$ $\frac{c_{k}e^{f_{i}+f_{j}}}{\sqrt{\xi_{jk}\xi_{ik}}e^{f_{i}+f_{j}}}\cdot\frac{1}{\sqrt{\xi_{jk}e^{2f_{k}}-2e^{f_{k}-f_{j}}}}\cdot\frac{1}{\sqrt{\xi_{ik}e^{2f_{k}}+2e^{f_{k}-f_{i}}}}\rightarrow+\infty.$
And $\widetilde{L}_{2}>0$. Then $\cosh\theta^{ij}_{k}=\widetilde{L}_{1}+\widetilde{L}_{2}\rightarrow+\infty$, which implies $\theta^{ij}_{k}\rightarrow+\infty$. The proof of the case that $\alpha_{i}=1$ is similar, so we omit it here.
For the case $\mathrm{(ii)}$, $\lim f_{i}=c,\lim f_{j}=+\infty,\ \lim f_{k}=-\infty$. Then $\lim\cosh l_{ij}:=\lim c^{\prime}_{k}e^{f_{j}}$$\lim\cosh l_{jk}=\lim(-1+\xi_{jk}e^{f_{j}+f_{k}})$ and
@@
where $c^{\prime}_{j},c^{\prime}_{k}$ are positive constants. If $\alpha_{i}=0$, then $\cosh l_{ik}\rightarrow 1$ and hence $l_{ik}\rightarrow 0$. Then $\cosh\theta^{ij}_{k}>\widetilde{L}_{2}>\frac{\cosh l_{ik}}{\sinh l_{ik}}\rightarrow+\infty$ and hence $\theta^{ij}_{k}\rightarrow+\infty$. If $\alpha_{i}=1$, similar to the proof of the case $\mathrm{(i)}$, we still have $\widetilde{L}_{1}\rightarrow+\infty$. This implies $\theta^{ij}_{k}\rightarrow+\infty$. This completes the proof. Q.E.D.
@@
Under the same assumptions as in Theorem 6.5. For any constants $c_{1},c_{2},c_{3}\in\mathbb{R}$, if one of the following conditions is satisfied
@@
$\lim f_{i}=-\infty,\ \lim f_{j}=+\infty,\ \lim f_{k}=+\infty$;
@@
$\lim f_{i}=-\infty,\ \lim f_{j}=+\infty,\ \lim f_{k}=-\infty$.
Then $\lim\theta^{jk}_{i}(f_{i},f_{j},f_{k})=+\infty$.
@@
Set $L_{1}=\frac{\cosh l_{jk}}{\sinh l_{ij}\sinh l_{ik}}$ and $L_{2}=\frac{\cosh l_{ij}\cosh l_{ik}}{\sinh l_{ij}\sinh l_{ik}}$, then $\cosh\theta^{jk}_{i}=L_{1}+L_{2}$. Since $\lim f_{i}=-\infty$, then (46) is reduced to
(57) $\cosh l_{ij}$ $\rightarrow\sqrt{1+\alpha_{j}e^{2f_{j}}}+\xi_{ij}e^{f_{i}+f_{j}},$
$\cosh l_{ik}$ $\rightarrow\sqrt{1+\alpha_{k}e^{2f_{k}}}+\xi_{ik}e^{f_{i}+f_{k}},$
$\cosh l_{jk}$ $=-\sqrt{(1+\alpha_{j}e^{2f_{j}})(1+\alpha_{k}e^{2f_{k}})}+\xi_{jk}e^{f_{j}+f_{k}}.$
For the case $\mathrm{(i)}$, $\lim f_{j}=+\infty,\ \lim f_{k}=+\infty$. By (57), we have $\lim\cosh l_{jk}:=\lim(c_{i}e^{f_{j}+f_{k}})$,
$\lim\cosh l_{ij}=\begin{cases}\lim(1+\xi_{ij}e^{f_{i}+f_{j}}),\ &\text{if}\ \alpha_{j}=0,\\ \lim(c_{k}e^{f_{j}}),\ &\text{if}\ \alpha_{j}=1,\end{cases}$
@@
$\lim\cosh l_{ik}=\begin{cases}\lim(1+\xi_{ik}e^{f_{i}+f_{k}}),\ &\text{if}\ \alpha_{k}=0,\\ \lim(c_{j}e^{f_{k}}),\ &\text{if}\ \alpha_{k}=1,\end{cases}$
@@
$L_{1}\rightarrow$ $\frac{c_{i}e^{f_{j}+f_{k}}}{\sqrt{(1+\xi_{ij}e^{f_{i}+f_{j}})^{2}-1}\cdot\sqrt{(1+\xi_{ik}e^{f_{i}+f_{k}})^{2}-1}}$
$=$ $\frac{c_{i}e^{f_{j}+f_{k}}}{\sqrt{\xi_{ij}\xi_{ik}}e^{f_{j}+f_{k}}}\cdot\frac{1}{\sqrt{\xi_{ij}e^{2f_{i}}+2e^{f_{i}-f_{j}}}}\cdot\frac{1}{\sqrt{\xi_{ik}e^{2f_{i}}+2e^{f_{i}-f_{k}}}}\rightarrow+\infty.$
And $L_{2}>0$. Thus $\cosh\theta^{jk}_{i}=L_{1}+L_{2}\rightarrow+\infty$, so $\theta^{jk}_{i}\rightarrow+\infty$. For the cases that $\alpha_{j}=0,\ \alpha_{k}=1$ and $\alpha_{j}=1,\ \alpha_{k}=0$, we still have $\theta^{jk}_{i}\rightarrow+\infty$ by similar proofs.
For the case $\mathrm{(ii)}$, $\lim f_{j}=+\infty,\ \lim f_{k}=c_{1}$. By (57), we have $\lim\cosh l_{jk}:=\lim(c^{\prime}_{i}e^{f_{j}})$ and
@@
where $c^{\prime}_{i},c^{\prime}_{j}$ are positive constants. If $\alpha_{k}=0$, then $l_{ik}\rightarrow 0$ and hence $L_{2}>\frac{\cosh l_{ik}}{\sinh l_{ik}}\rightarrow+\infty$. This implies $\cosh\theta^{jk}_{i}\rightarrow+\infty$. If $\alpha_{k}=1$, then by the assumption, $\alpha_{j}=0$. Hence, $\lim\cosh l_{ij}=\lim(1+\xi_{ij}e^{f_{i}+f_{j}})$, $\lim\cosh l_{jk}=\lim(c^{\prime}_{i}e^{f_{j}})$ and $\lim\cosh l_{ik}=c^{\prime}_{j}$. Then
$L_{1}\rightarrow$ $\frac{c^{\prime}_{i}e^{f_{j}}}{\sqrt{(1+\xi_{ij}e^{f_{i}+f_{j}})^{2}-1}\cdot\sqrt{(c^{\prime}_{j})^{2}-1}}$
$=$ $\frac{c^{\prime}_{i}e^{f_{j}}}{\sqrt{\xi_{ij}}\sqrt{(c^{\prime}_{j})^{2}-1}e^{f_{j}}}\cdot\frac{1}{\sqrt{\xi_{ij}e^{2f_{i}}+2e^{f_{i}-f_{j}}}}\rightarrow+\infty.$
And $L_{2}>0$. Thus $\cosh\theta^{jk}_{i}=L_{1}+L_{2}\rightarrow+\infty$, so $\theta^{jk}_{i}\rightarrow+\infty$.
For the case $\mathrm{(iii)}$, $\lim f_{j}=c_{2},\ \lim f_{k}=c_{3}$. By (57), we have $\lim\cosh l_{jk}:=c^{\prime\prime}_{i}$ and
@@
where $c^{\prime\prime}_{i},c^{\prime\prime}_{k}$ are positive constants. If $\alpha_{j}=0$, then $l_{ij}\rightarrow 0$ and hence $L_{2}\rightarrow+\infty$. This implies $\theta^{jk}_{i}\rightarrow+\infty$. If $\alpha_{j}=1$, then by the assumption, $\alpha_{k}=0$. According to the arguments in the case $\mathrm{(ii)}$, we still have $\theta^{jk}_{i}\rightarrow+\infty$.
For the case $\mathrm{(iv)}$, $\lim f_{j}=+\infty,\ \lim f_{k}=-\infty$. By (57), we have $\lim\cosh l_{ik}=1$ and hence $l_{ik}\rightarrow 0$. Then $\cosh\theta^{jk}_{i}>L_{2}\geq\frac{\cosh l_{ik}}{\sinh l_{ik}}\rightarrow+\infty$ and then $\theta^{jk}_{i}\rightarrow+\infty$. This completes the proof. Q.E.D.
@@
According to the proof of the case $\mathrm{(iii)}$ in Lemma 6.8, if $\alpha_{j}=\alpha_{k}=1$, $\lim\cosh l_{ij}=c^{\prime\prime}_{k}$, $\lim\cosh l_{jk}=c^{\prime\prime}_{i}$ and $\lim\cosh l_{ik}=c^{\prime}_{j}$, then $\cosh\theta^{jk}_{i}$, $\cosh\theta^{ik}_{j}$ and $\cosh\theta^{ij}_{k}$ all tend to finite constants, which are far away from 1 and +$\infty$+ ∞.
@@
Under the same assumptions as in Theorem 6.5. For any constants $c_{1},c_{2},c_{3}\in\mathbb{R}$, if one of the following conditions is satisfied
@@
$\lim f_{j}=+\infty,\ \lim f_{i}=c_{2},\ \lim f_{k}=c_{3}$.
Then $\lim\theta^{ik}_{j}(f_{i},f_{j},f_{k})=0$.
@@
For the case $\mathrm{(i)}$, by (56), we have $\lim\cosh l_{ij}:=\lim(c_{k}e^{f_{j}})$, $\lim\cosh l_{ik}:=c_{j}e^{f_{k}}$ and $\lim\cosh l_{jk}:=\lim(c_{i}e^{f_{j}+f_{k}})$, where $c_{i},c_{j},c_{k}$ are positive constants. Then
@@
which implies $\theta^{ik}_{j}\rightarrow 0$. The proof of the case $\mathrm{(ii)}$ is similar, so we omit it here. Q.E.D.
Proof of Theorem 6.5: Let the image of $K$ is $X$. The openness of $X$ in >$\mathbb{R}^{N}_{>0}$ follows from the injectivity of the map $f$. To prove the closeness of $X$ in >$\mathbb{R}^{N}_{>0}$, we prove that if a sequence $f^{(m)}$ in $\Omega(\xi)=\bigcap_{\{ijk\}\in F}\Omega_{ijk}(\xi)$ satisfies $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$, then there exists a subsequence, say $f^{(m)}$, such that $\lim_{m\rightarrow+\infty}f^{(m)}$ is in $\Omega(\xi)$.
Suppose otherwise, there exists a subsequence $f^{(m)}$, such that $\lim_{m\rightarrow+\infty}f^{(m)}$ is in the boundary of $\Omega(\xi)$. The boundary of the admissible space $\Omega(\xi)$ in $[-\infty,+\infty]^{N}$ consists of the following three parts:
@@
For the case $\mathrm{(i)}$, in a right-angled hyperbolic hexagon $\{ijk\}\in F$ with lengths $l_{ij},l_{jk},l_{ik}$ and opposite lengths of boundaries $\theta^{ij}_{k},\theta^{jk}_{i},\theta^{ik}_{j}$. Without loss of generality, we assume $l_{ij}\rightarrow 0$. By the cosine law,
which contradicts the assumption that $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$.
We will prove that the above cases $\mathrm{(ii)}$, $\mathrm{(iii)}$ and $\mathrm{(iv)}$ can not hold on $\mathcal{U}^{(n)}_{ijk}(\xi)$, $n=1,2,4,5.$
(1): For $(0,0,0)$ type admissible space $\mathcal{U}^{(1)}_{ijk}(\xi)$ in (48), we have $u_{j}+u_{k}>\log\frac{2}{\xi_{jk}}$. By Remark 2.23, $u_{j}=f_{j}$ and $u_{k}=f_{k}$, then $f_{j}+f_{k}$ has a lower bound. Hence, $f_{j},\ f_{k}$ can not tend to $\infty$- ∞ simultaneously.
@@
(2): For $(1,0,0)$ type admissible space $\mathcal{U}^{(2)}_{ijk}(\xi)$ in (49), we have $u_{j}+u_{k}>\log\frac{2}{\xi_{jk}}$. The proof is the same as that of $(0,0,0)$ type admissible space $\mathcal{U}^{(1)}_{ijk}(\xi)$, so we omit it here.
(3): For $(0,0,1)$ type admissible space $\mathcal{U}^{(4)}_{ijk}(\xi)$ in (50), the positions of $j$ and $k$ are not interchangeable.
(i):
In the admissible space (51), we have $u_{j}+u_{k}>\log\frac{1}{\xi_{jk}}$. By (47) and Remark 2.23, $u_{j}=f_{j}\in\mathbb{R}$ and $u_{k}=-\mathrm{arcsinh}(e^{-f_{k}})\in\mathbb{R}_{0}$, then $f_{j}$ can not tend to $\infty$- ∞.
@@
If $f_{j}^{(m)}\rightarrow+\infty$ and $f_{k}^{(m)}\rightarrow-\infty$, by Lemma 6.7 (ii), then $(\theta^{ij}_{k})^{(m)}\rightarrow+\infty$ and hence $K^{(m)}_{k}\rightarrow+\infty$. If $f_{j}^{(m)}\rightarrow+\infty$ and $f_{k}^{(m)}\nrightarrow-\infty$, by Lemma 6.10, then $(\theta^{ik}_{j})^{(m)}\rightarrow 0$ and hence $K^{(m)}_{j}\rightarrow 0$. Hence, the subsequence $f^{(m)}_{j}$ can not tend to +$\infty$+ ∞ in the admissible space (51).
@@
(ii):
In the admissible space (52), we have $u_{j}+u_{k}>\log\frac{1}{\xi_{jk}},\ u_{i}+u_{k}>\log(-\xi_{ik})$. According to the case $\mathrm{(i)}$, $f_{j}$ can not tend to $\infty$- ∞. Similarly, combining (47) and Remark 2.23 gives $u_{i}=-f_{i}$. Then $f_{i}$ can not tend to +$\infty$+ ∞.
If $f_{i}^{(m)}\rightarrow-\infty$, by Lemma 6.8, then $(\theta^{jk}_{i})^{(m)}\rightarrow+\infty$ and hence $K^{(m)}_{i}\rightarrow+\infty$. Hence, the subsequence $f^{(m)}_{i}$ can not tend to $\infty$- ∞ in the admissible space (52).
If $f_{j}^{(m)}\rightarrow+\infty$ and $f_{k}^{(m)}\rightarrow-\infty$, then $f_{i}^{(m)}\rightarrow-\infty$ to ensure $u_{i}+u_{k}>\log(-\xi_{ik})$. This case has been included in the above arguments. If $f_{j}^{(m)}\rightarrow+\infty$ and $f_{k}^{(m)}\nrightarrow-\infty$, by Lemma 6.10, then $(\theta^{ik}_{j})^{(m)}\rightarrow 0$ and hence $K^{(m)}_{j}\rightarrow 0$. Hence, the subsequence $f^{(m)}_{j}$ can not tend to +$\infty$+ ∞ in the admissible space (52).
@@
(4): For $(1,0,1)$ type admissible space $\mathcal{U}^{(5)}_{ijk}(\xi)$ in (53), we have $u_{j}+u_{k}>\log\frac{1}{\xi_{jk}}$. The proof is the same as that of $(0,0,1)$ type admissible space (51), so we omit it here. Q.E.D.
Remark 6.11.
For $(0,1,1)$ type admissible space $\mathcal{U}^{(10)}_{ijk}(\xi)$ in (54) and $(1,1,1)$ type admissible space $\mathcal{U}^{(11)}_{ijk}(\xi)$ in (55), we are unable to prove the image of $K$ is >$\mathbb{R}^{N}_{>0}$. The reasons follow from Remark 6.9.

### v3

@@ 8.4. Existence of the mixed discrete conformal structure I
$\widetilde{L}_{1}\rightarrow$ $\frac{c_{k}e^{f_{i}+f_{j}}}{\sqrt{(-1+\eta_{jk}e^{f_{j}+f_{k}})^{2}-1}\cdot\sqrt{(1+\eta_{ik}e^{f_{i}+f_{k}})^{2}-1}}$
$=$ $\frac{c_{k}e^{f_{i}+f_{j}}}{\sqrt{\eta_{jk}\eta_{ik}}e^{f_{i}+f_{j}}}\cdot\frac{1}{\sqrt{\eta_{jk}e^{2f_{k}}-2e^{f_{k}-f_{j}}}}\cdot\frac{1}{\sqrt{\eta_{ik}e^{2f_{k}}+2e^{f_{k}-f_{i}}}}\rightarrow+\infty.$
Since $\widetilde{L}_{2}>0$, it follows that $\cosh\theta^{ij}_{k}=\widetilde{L}_{1}+\widetilde{L}_{2}\rightarrow+\infty$, which implies $\theta^{ij}_{k}\rightarrow+\infty$. The proof of the case that $\alpha_{i}=1$ is similar, so we omit it here.
For the case (ii), it follows that $\lim\cosh l_{ij}:=\lim c^{\prime}_{k}e^{f_{j}}$, $\lim\cosh l_{jk}=\lim(-1+\eta_{jk}e^{f_{j}+f_{k}})$, and
@@
where $c^{\prime}_{j},c^{\prime}_{k}$ are positive constants. If $\alpha_{i}=0$, then $\cosh l_{ik}\rightarrow 1$, and hence $l_{ik}\rightarrow 0$. This implies $\cosh\theta^{ij}_{k}>\widetilde{L}_{2}>\frac{\cosh l_{ik}}{\sinh l_{ik}}\rightarrow+\infty$, and hence $\theta^{ij}_{k}\rightarrow+\infty$. If $\alpha_{i}=1$, similar to the proof of the case (i), we derive $\widetilde{L}_{1}\rightarrow+\infty$. Consequently, $\theta^{ij}_{k}\rightarrow+\infty$. This completes the proof.
Q.E.D.
@@
Under the same assumptions as those in Lemma 8.7, if one of the following conditions is satisfied
@@
$\lim f_{i}=-\infty,\ \lim f_{j}=+\infty,\ \lim f_{k}=+\infty$,
@@
$\lim f_{i}=-\infty,\ \lim f_{j}=+\infty,\ \lim f_{k}=-\infty$,
where $c_{1},c_{2},c_{3}\in\mathbb{R}$ are constants. Then $\lim\theta^{jk}_{i}(f_{i},f_{j},f_{k})=+\infty$.
@@
Set $L_{1}=\frac{\cosh l_{jk}}{\sinh l_{ij}\sinh l_{ik}}$ and $L_{2}=\frac{\cosh l_{ij}\cosh l_{ik}}{\sinh l_{ij}\sinh l_{ik}}$, then $\cosh\theta^{jk}_{i}=L_{1}+L_{2}$. Given that $\lim f_{i}=-\infty$, the formula (53) reduces to
(64) $\cosh l_{ij}$ $\rightarrow\sqrt{1+\alpha_{j}e^{2f_{j}}}+\eta_{ij}e^{f_{i}+f_{j}},$
$\cosh l_{ik}$ $\rightarrow\sqrt{1+\alpha_{k}e^{2f_{k}}}+\eta_{ik}e^{f_{i}+f_{k}},$
$\cosh l_{jk}$ $=-\sqrt{(1+\alpha_{j}e^{2f_{j}})(1+\alpha_{k}e^{2f_{k}})}+\eta_{jk}e^{f_{j}+f_{k}}.$
For the case (i), it follows from (64) that $\lim\cosh l_{jk}:=\lim(c_{i}e^{f_{j}+f_{k}})$,
$\lim\cosh l_{ij}=\begin{cases}\lim(1+\eta_{ij}e^{f_{i}+f_{j}}),\ &\text{if}\ \alpha_{j}=0,\\ \lim(c_{k}e^{f_{j}}),\ &\text{if}\ \alpha_{j}=1,\end{cases}$
@@
$\lim\cosh l_{ik}=\begin{cases}\lim(1+\eta_{ik}e^{f_{i}+f_{k}}),\ &\text{if}\ \alpha_{k}=0,\\ \lim(c_{j}e^{f_{k}}),\ &\text{if}\ \alpha_{k}=1,\end{cases}$
@@
$L_{1}\rightarrow$ $\frac{c_{i}e^{f_{j}+f_{k}}}{\sqrt{(1+\eta_{ij}e^{f_{i}+f_{j}})^{2}-1}\cdot\sqrt{(1+\eta_{ik}e^{f_{i}+f_{k}})^{2}-1}}$
$=$ $\frac{c_{i}e^{f_{j}+f_{k}}}{\sqrt{\eta_{ij}\eta_{ik}}e^{f_{j}+f_{k}}}\cdot\frac{1}{\sqrt{\eta_{ij}e^{2f_{i}}+2e^{f_{i}-f_{j}}}}\cdot\frac{1}{\sqrt{\eta_{ik}e^{2f_{i}}+2e^{f_{i}-f_{k}}}}\rightarrow+\infty.$
Since $L_{2}>0$, it follows that $\cosh\theta^{jk}_{i}=L_{1}+L_{2}\rightarrow+\infty$, which implies $\theta^{jk}_{i}\rightarrow+\infty$. For the cases that $\alpha_{j}=0,\ \alpha_{k}=1$ and $\alpha_{j}=1,\ \alpha_{k}=0$, analogous reasoning yields $\theta^{jk}_{i}\rightarrow+\infty$.
For the case (ii), it follows from (64) that $\lim\cosh l_{jk}:=\lim(c^{\prime}_{i}e^{f_{j}})$, and
@@
where $c^{\prime}_{i},c^{\prime}_{j}$ are positive constants. If $\alpha_{k}=0$, then $l_{ik}\rightarrow 0$, and hence $L_{2}>\frac{\cosh l_{ik}}{\sinh l_{ik}}\rightarrow+\infty$. This implies $\cosh\theta^{jk}_{i}\rightarrow+\infty$. If $\alpha_{k}=1$, then by the assumption, we have $\alpha_{j}=0$. Hence, $\lim\cosh l_{ij}=\lim(1+\eta_{ij}e^{f_{i}+f_{j}})$, $\lim\cosh l_{jk}=\lim(c^{\prime}_{i}e^{f_{j}})$, and $\lim\cosh l_{ik}=c^{\prime}_{j}$. Consequently,
$L_{1}\rightarrow$ $\frac{c^{\prime}_{i}e^{f_{j}}}{\sqrt{(1+\eta_{ij}e^{f_{i}+f_{j}})^{2}-1}\cdot\sqrt{(c^{\prime}_{j})^{2}-1}}$
$=$ $\frac{c^{\prime}_{i}e^{f_{j}}}{\sqrt{\eta_{ij}}\sqrt{(c^{\prime}_{j})^{2}-1}e^{f_{j}}}\cdot\frac{1}{\sqrt{\eta_{ij}e^{2f_{i}}+2e^{f_{i}-f_{j}}}}\rightarrow+\infty.$
Since $L_{2}>0$, it follows that $\cosh\theta^{jk}_{i}=L_{1}+L_{2}\rightarrow+\infty$, which implies $\theta^{jk}_{i}\rightarrow+\infty$.
For the case (iii), it follows from (64) that $\lim\cosh l_{jk}:=c^{\prime\prime}_{i}$, and
@@
where $c^{\prime\prime}_{i},c^{\prime\prime}_{k}$ are positive constants. If $\alpha_{j}=0$, then $l_{ij}\rightarrow 0$, and hence $L_{2}\rightarrow+\infty$. This implies $\theta^{jk}_{i}\rightarrow+\infty$. If $\alpha_{j}=1$, then by the assumption, $\alpha_{k}=0$. According to the arguments in the case (ii), we still have $\theta^{jk}_{i}\rightarrow+\infty$.
For the case (iv), it follows from (64) that $\lim\cosh l_{ik}=1$, and hence $l_{ik}\rightarrow 0$. Consequently, $\cosh\theta^{jk}_{i}>L_{2}\geq\frac{\cosh l_{ik}}{\sinh l_{ik}}\rightarrow+\infty$, which implies $\theta^{jk}_{i}\rightarrow+\infty$. This completes the proof. Q.E.D.
@@
As shown in the proof of case (iii) in Lemma 8.9, if $\alpha_{j}=\alpha_{k}=1$, then $\lim\cosh l_{ij}=c^{\prime\prime}_{k}$, $\lim\cosh l_{jk}=c^{\prime\prime}_{i}$, and $\lim\cosh l_{ik}=c^{\prime}_{j}$. In this case, $\cosh\theta^{jk}_{i}$, $\cosh\theta^{ik}_{j}$ and $\cosh\theta^{ij}_{k}$ all tend to finite constants, which neither approach 1 nor +$\infty$+ ∞. Consequently, no contradiction can be derived from this case in the subsequent analysis, thereby justifying the exclusion of the case that $\alpha_{j}=\alpha_{k}=1$.
@@
Under the same assumptions as those in Lemma 8.7, if one of the following conditions is satisfied
@@
$\lim f_{j}=+\infty,\ \lim f_{i}=c_{2},\ \lim f_{k}=c_{3}$,
where $c_{1},c_{2},c_{3}\in\mathbb{R}$ are constants. Then $\lim\theta^{ik}_{j}(f_{i},f_{j},f_{k})=0$.
@@
For the case (i), by (63), we have $\lim\cosh l_{ij}:=\lim(c_{k}e^{f_{j}})$, $\lim\cosh l_{ik}:=c_{j}e^{f_{k}}$, and $\lim\cosh l_{jk}:=\lim(c_{i}e^{f_{j}+f_{k}})$, where $c_{i},c_{j},c_{k}$ are positive constants. Then
@@
which implies $\theta^{ik}_{j}\rightarrow 0$. The proof of the case (ii) is similar, so we omit it here. Q.E.D.
Proof of Theorem 8.6: Let $X$ be the image of $K$. The openness of $X$ in >$\mathbb{R}^{N}_{>0}$ follows from the smooth injectivity of the map. To prove the closedness of $X$ in >$\mathbb{R}^{N}_{>0}$, we prove that if a sequence $\{f^{(m)}\}\in\Omega(\eta)=\bigcap_{\{ijk\}\in F}\Omega_{ijk}(\eta)$ satisfies $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$, then there exists a subsequence, still denoted by $\{f^{(m)}\}$, such that $\lim_{m\rightarrow+\infty}f^{(m)}\in\Omega(\eta)$.
Suppose otherwise, there exists a subsequence $\{f^{(m)}\}$, such that $\lim_{m\rightarrow+\infty}f^{(m)}$ lies on the boundary of $\Omega(\eta)$. The boundary of the admissible space $\Omega(\eta)$ in $[-\infty,+\infty]^{N}$ consists of the following five parts:
@@
(v):
The proof for case (i) is analogous to that of Theorem 3.5, and is thus omitted here.
We first establish that the aforementioned cases (ii), (iii), and (iv) cannot hold on $\mathcal{U}^{(n)}(\eta)=\bigcap_{\{ijk\}\in F}\mathcal{U}^{(n)}_{ijk}(\eta)$, where $n=1,2,4,5$. The corresponding results are stated as follows.
  • (1)
    For $(0,0,0)$-type admissible space $\mathcal{U}^{(1)}_{ijk}(\eta)$ in (55), we have $u_{j}+u_{k}>\log\frac{2}{\eta_{jk}}$. By Remark 2.22, $u_{j}=f_{j}$ and $u_{k}=f_{k}$, then $f_{j}+f_{k}$ has a lower bound. Hence, both $f_{j}$ and $f_{k}$ cannot tend to $\infty$- ∞ simultaneously.
@@
  • (2)
    For $(1,0,0)$-type admissible space $\mathcal{U}^{(2)}_{ijk}(\eta)$ in (56), we have $u_{j}+u_{k}>\log\frac{2}{\eta_{jk}}$. The proof follows identically to that of the $(0,0,0)$-type admissible space $\mathcal{U}^{(1)}_{ijk}(\eta)$, and thus is omitted here.
  • (3)
    For $(0,0,1)$-type admissible space $\mathcal{U}^{(4)}_{ijk}(\eta)$ in (57), the positions of $j$ and $k$ are not interchangeable.
    • (a)
      In the admissible space (58), we have $u_{j}+u_{k}>\log\frac{1}{\eta_{jk}}$. Combining (54) and Remark 2.22 yields $u_{j}=f_{j}\in\mathbb{R}$ and $u_{k}=-\mathrm{arcsinh}\,(e^{-f_{k}})\in\mathbb{R}_{<0}$. It follows that $f_{j}$ cannot tend to $\infty$- ∞.
@@
      If $f_{j}^{(m)}\rightarrow+\infty$ and $f_{k}^{(m)}\rightarrow-\infty$, by Lemma 8.8 (ii), then $(\theta^{ij}_{k})^{(m)}\rightarrow+\infty$ and hence $K^{(m)}_{k}\rightarrow+\infty$. If $f_{j}^{(m)}\rightarrow+\infty$ and $f_{k}^{(m)}\nrightarrow-\infty$, by Lemma 8.11, then $(\theta^{ik}_{j})^{(m)}\rightarrow 0$ and hence $K^{(m)}_{j}\rightarrow 0$. Consequently, the subsequence $f^{(m)}_{j}$ cannot tend to +$\infty$+ ∞ in the admissible space (58).
@@
    • (b)
      In the admissible space (59), we have $u_{j}+u_{k}>\log\frac{1}{\eta_{jk}},\ u_{i}+u_{k}>\log(-\eta_{ik})$. The case (a) shows that $f_{j}$ cannot tend to $\infty$- ∞. Similarly, combining (54) and Remark 2.22 gives $u_{i}=-f_{i}$. Thus $f_{i}$ cannot tend to +$\infty$+ ∞.
      If $f_{i}^{(m)}\rightarrow-\infty$, by Lemma 8.9, then $(\theta^{jk}_{i})^{(m)}\rightarrow+\infty$ and hence $K^{(m)}_{i}\rightarrow+\infty$. Consequently, the subsequence $f^{(m)}_{i}$ cannot tend to $\infty$- ∞ in the admissible space (59).
      If $f_{j}^{(m)}\rightarrow+\infty$ and $f_{k}^{(m)}\rightarrow-\infty$, then $f_{i}^{(m)}\rightarrow-\infty$ must hold to satisfy the inequality $u_{i}+u_{k}>\log(-\eta_{ik})$. This case has been included in the above arguments. If $f_{j}^{(m)}\rightarrow+\infty$ and $f_{k}^{(m)}\nrightarrow-\infty$, by Lemma 8.11, then $(\theta^{ik}_{j})^{(m)}\rightarrow 0$ and hence $K^{(m)}_{j}\rightarrow 0$. Consequently, the subsequence $f^{(m)}_{j}$ cannot tend to +$\infty$+ ∞ in the admissible space (59).
@@
  • (4)
    For $(1,0,1)$-type admissible space $\mathcal{U}^{(5)}_{ijk}(\eta)$ in (60), we have $u_{j}+u_{k}>\log\frac{1}{\eta_{jk}}$. The proof follows identically to that of the $(0,0,1)$-type admissible space (58), and thus is omitted here.
From the foregoing proofs, it follows that the generalized angle tends to either 00 or +$\infty$+ ∞ on the admissible space $\mathcal{U}^{(n)}(\eta)$, which is composed of the single admissible subspace $\mathcal{U}^{(n)}_{ijk}(\eta)$ for $n=1,2,4,5$. Furthermore, for the case (v), the proof of Theorem 3.5 demonstrates that the generalized angle $\theta_{j}^{ik}$ tends to either 00 or +$\infty$+ ∞ on $\mathcal{U}^{(0)}_{ijk}(\eta)$. Through a lengthy yet analogous analysis, one can establish that the generalized angle tends to either 00 or +$\infty$+ ∞ on $\widetilde{\mathcal{U}}(\eta)$, which is formed by mixed admissible subspaces $\mathcal{U}^{(n)}_{ijk}(\eta)$ with $n=0,1,2,4,5$. Accordingly, the aforementioned cases (ii), (iii), (iv), and (v) cannot hold on $\widetilde{\mathcal{U}}(\eta)$. This completes the proof.
Q.E.D.